Please help - I have such a hard time with word problems!
The income from a production was $11,400. The price of an advanced ticket was $3 and at the door tickets were $5. If three thousand tickets were sold, how many of each kind were sold?

Found 2 solutions by funmath, stanbon:
Answer by funmath(2873)

The income from a production was $11,400. The price of an advanced ticket was $3 and at the door tickets were $5. If three thousand tickets were sold, how many of each kind were sold?
Let advanced tickets=x
Let door tickets =y
There were 3000 tickets all together, so x+y=3000
x+y=3000
-x+x+y=3000-x
y=3000-x
Therefore
advanced tickets are: x
Door tickets are: 3000-x
You get $3 per advanced ticket: 3x
You get $5 per door ticket: 5(3000-x)
You got $11,400 altogether, so 3x+5(3000-x)=11400
3x+5(3000)+5(-x)=11400
3x+15000-5x=11400
-2x+15000=11400
-2x+15000-15000=11400-15000
-2x=-3600
-2x/-2=-3600/-2
x=1800
Advance tickets: x=1800
Door tickets: 3000-x=3000-1800=1200
Happy Calculating!!!!

Answer by stanbon(48564)

The income from a production was $11,400. The price of an advanced ticket was $3 and at the door tickets were $5. If three thousand tickets were sold, how many of each kind were sold?
--------------
Let the number of advanced tickets sold be "x" ; Value of these is 3x dollars
Number of door tickets sold is "3000-x" ; Value of these is 5(3000-x)=15000-5x
dollars.
---------------
EQUATION:
value + value = 11400 dollars
3x + 15000-5x = 11400
-2x = -3600
x=1800 (# of advanced tickets sold)
3000-x = 3000-1800 = 1200 (# of door tickets sold)
=============
Cheers,
Stan H.